OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1 A. -69x + 10y = -51 B. 69x + 10y = 51 C. -69x + 10y = 51 D. 69x + 10y = -51

OpenStudy (anonymous):

Ax+By=C --> standard form so rearrange your given equation into standard form first then I'll tell you what you need to do next :)

OpenStudy (anonymous):

I'm really not sure how to do this problem...

OpenStudy (anonymous):

right now your equation is in the form y=mx+b change it to Ax+By=C so y = -6.9x + 5.1 will be 6.9x+y=5.1 by adding both sides by 6.9x then your question says " integer coefficients." which means that all the coefficients should be whole numbers.. do this by multiplying each term by 10 therefore you will have : 69x+10y=51 is it clear?

OpenStudy (anonymous):

sorry late repz.

OpenStudy (anonymous):

Thank you!